Perfectly Competitive firm faces a market price of birr 40 and has the following

Cost function: STC = 5800+ 20Q+0.02Q2.

A. What quantity of output is best for this firm in the short-run? Why?

B. Should firm attempt to change some price other than the market price of 40? Why or why not?


1
Expert's answer
2021-12-20T19:48:10-0500

(a) Given that:

"STC=5800+20Q+0.02Q^{2}"

"MC=20+0.04Q"

For maximum profit MC=P

But "p=40"

Therefore, "40=20+0.04Q"

"40-20=0.04Q"

"20=0.04Q"

"Q=\frac{20}{0.04}=500"

Quantity of output best for this firm is 500


(b) Total revenue is given by:

"TR=P\times Q"

"TR=40\times 500=20000"

Total cost is given by:

"TC=5800+20Q +0.02Q^{2}"

"TC=5800+20\times 500 +0.02\times (0.02\times 500^{2})=20800"

Profits is equals to Total revenue minus total cost.

"Profits=20000-20800=-800"

The firm is running at a loss.

The firm should attempt to change some price other than the market price of 40 to help it earn some profit instead of making losses.